OpenStudy (anonymous):

PLEASE HELP BIG GRADE Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (love_ranaa):

Answers: f(3) = f(2) + f(1) => 3 + 2 = 5 f(4) = f(3) + f(2) => 5 + 3 = 8 f(5) = f(4) + f(3) => 8 + 5 = 13 Therefore the sequence is 2, 3, 5, 8, 13,.....a Fibonacci sequence. :)

OpenStudy (anonymous):

@Love_Ranaa that is not in the choices though?

OpenStudy (love_ranaa):

Oh Im sorry. Wrong numbers ..

OpenStudy (welshfella):

f(2) + f(1) = 2+1 = 3

OpenStudy (love_ranaa):

Its C

OpenStudy (welshfella):

so f(3) = 3

OpenStudy (welshfella):

after the second term each term is obtained by adding the previous 2 terms